buddyboy:

m∠BAD = 129° m∠BAC = (2x −1)° m∠CAD = (3x + 5)° Using the angle addition postulate, write and solve an equation for x. I don't want the answer I just want to know how to answer it for lack of Beter words I need help completing

elizabeth22323:

BAC AND CAD

Convert:

@elizabeth22323 wrote:
BAC AND CAD
thats crazy how u did the one thing they said not to do

buddyboy:

wdym? it needs me to solve for X

elizabeth22323:

@convert wrote:
@elizabeth22323 wrote:
BAC AND CAD
thats crazy how u did the one thing they said not to do
Sorry

Convert:

@buddyboy wrote:
wdym? it needs me to solve for X
he said i dont want the answer i just wanna know how to answer it

elizabeth22323:

@buddyboy wrote:
wdym? it needs me to solve for X
equal them to 360 or 180

buddyboy:

so the answer is just BAC and CAD

elizabeth22323:

@buddyboy wrote:
so the answer is just BAC and CAD
no take both equations and equal them to 129

Convert:

If angles BAC and CAD are adjacent you would equal them to the measure of angle BAD which is what the angle addition postulate is

buddyboy:

@convert wrote:
If angles BAC and CAD are adjacent you would equal them to the measure of angle BAD which is what the angle addition postulate is
ok but I dont know how to do that

Convert:

So setting up the problem would made it m∠BAC+m∠CAD = m∠CAD or when you plug the numbers in: (2x −1)+(3x+5) = 129 degrees

Convert:

@buddyboy wrote:
@convert wrote:
If angles BAC and CAD are adjacent you would equal them to the measure of angle BAD which is what the angle addition postulate is
ok but I dont know how to do that
which part are you having trouble with?

buddyboy:

@convert wrote:
@buddyboy wrote:
@convert wrote:
If angles BAC and CAD are adjacent you would equal them to the measure of angle BAD which is what the angle addition postulate is
ok but I dont know how to do that
which part are you having trouble with?
i honestly dont know I have been up all day doing school work an I have been stressed and this is making it worst

buddyboy:

@buddyboy wrote:
@convert wrote:
as in the question n ot you
@buddyboy wrote:
@convert wrote:
If angles BAC and CAD are adjacent you would equal them to the measure of angle BAD which is what the angle addition postulate is
ok but I dont know how to do that
which part are you having trouble with?
i honestly dont know I have been up all day doing school work an I have been stressed and this is making it worst

Convert:

 So when setting it up it would look like that

buddyboy:

yes do you want me to send a picture of it

Convert:

The angle of BAD is equal to 129 degrees so the big angle

Convert:

no its fine does it look like what i just drew?

buddyboy:

yes

Convert:

Ok so do you understand which angles are which? Like which angle is equal to 129 degrees and which is 2x-1 or 3x+5?

buddyboy:

yes i think i do

Convert:

Ok so which one would be the biggest angle do you think?

Convert:

the biggest one you can create off of what the angles you have right now

Convert:

ignore the what

buddyboy:

so the big angle(both of the smaller angles together) is 129

Convert:

exactly and the small angles are BAC = (2x −1)° and CAD = (3x + 5)°

Convert:

so if both the smaller angles make up the big angle you would set them to equal each other

Convert:

Kind of like splitting up the number 2 into 1 and 1

buddyboy:

so if i find x the i can find how many deg both the other angles are, right?

Convert:

two 1's make up a 2 if that makes sense

Convert:

yup the x in this case would be degrees so you would solve for x then plug it back in to each angle to get how many degrees each angle is equal to

Convert:

Try it out for me and let me know what you get while walking me through the steps

buddyboy:

I am soo sorry i dont know where to start

Convert:

@buddyboy wrote:
I am soo sorry i dont know where to start
its ok! ill start it off for you the equation would be (2x-1)+(3x-5) = 129

Convert:

step 1 would be combine like terms where we can combine the x values and the regular numbers

buddyboy:

ok so i would combine like terms like 2x and 3x giving me 5x

Convert:

perfect! yes you would

Convert:

then the other 2?

buddyboy:

and then will i add1 to -5?

Convert:

careful the - is still in front of the 1

Convert:

and the 5 isn't negative

buddyboy:

ok i thought you did the invers

Convert:

@convert wrote:
@buddyboy wrote:
I am soo sorry i dont know where to start
its ok! ill start it off for you the equation would be (2x-1)+(3x-5) = 129
just noticed i got it wrong on here sorry

buddyboy:

its not a -5